# Beyond Browsing: API-Based Web Agents

### v1

@@ Abstract
Web browsers are a portal to the internet, where much of human activity is undertaken. Thus, there has been significant research work in AI agents that interact with the internet through web browsing. However, there is also another interface designed specifically for machine interaction with online content: application programming interfaces (APIs). In this paper we ask – what if we were to take tasks traditionally tackled by browsing agents, and give AI agents access to APIs? To do so, we propose two varieties of agents: (1) an API-calling agent that attempts to perform online tasks through APIs only, similar to traditional coding agents, and (2) a Hybrid Agent that can interact with online data through both web browsing and APIs. In experiments on WebArena, a widely-used and realistic benchmark for web navigation tasks, we find that API-based agents outperform web browsing agents. Hybrid Agents out-perform both others nearly uniformly across tasks, resulting in a more than 20.0% absolute improvement over web browsing alone, achieving a success rate of 35.8%, achiving the SOTA performance among task-agnostic agents. These results strongly suggest that when APIs are available, they present an attractive alternative to relying on web browsing alone.111Code is available at https://github.com/yueqis/API-Based-Agent; Project Webpage is available at https://yueqis.github.io/API-Based-Agent/
@@ 1 Introduction
Web agents use browsers as a interface to facilitate humans in performing daily tasks such as online shopping, online planning, trip planning, and other work-related tasks (Liu et al., 2018; Li et al., 2020; Rawles et al., 2023; Patil et al., 2023; Pan et al., 2024; Chen et al., 2024a; Huang et al., 2024; Durante et al., 2024). Existing web agents typically operate within the space of graphical user interfaces (GUI) (Zhang et al., 2023; Zhou et al., 2023; Zheng et al., 2024), using action spaces that simulate human-like keyboard and mouse operations, such as clicking and typing. To observe web pages, common approaches include using accessibility trees, a simplified version of the HTML DOM tree, as the input to text-based models (Zhou et al., 2023; Drouin et al., 2024a), or multi-modal, screenshot-based models (Koh et al., 2024a; Xie et al., 2024; You et al., 2024; Hong et al., 2023). However, regardless of the method of interaction with web sites, there is no getting around the fact that these sites were originally designed for human consumption, and may not be the ideal interface for machines.
@@
Notably, there is another interface designed specifically for machine interaction with online content: application programming interfaces (APIs). APIs allow machines to communicate directly with the backend of a web service (Branavan et al., 2009), sending and receiving data in machine-friendly formats such as JSON or XML (Meng et al., 2018; Xu et al., 2021). Nonetheless, whether AI agents can effectively use APIs to tackle real-world online tasks, and the conditions under which this is possible, remain unstudied in the scientific literature. In this work, we explore methods for tackling tasks normally framed as web-navigation tasks with an expanded action space to interact with APIs. To do so, we develop new API-based agents that directly interact with web services via API calls, as depicted in Figure 1. This method bypasses the need to interact with web page GUIs through simulated clicks.
@@
At the same time, not all websites have extensive API support, in which case web browsing actions may still be required. To address these cases, we explore a hybrid approach that combines API-based agents with web-browsing agents, as described in Figure 1. By implementing an agent capable of interleaving API calls and web browsing, we found that agents benefit from the flexibility of this hybrid model. When APIs are available and well-documented, the agent can directly interact with the web services. For websites with limited API support, the agent seamlessly switches to web browsing mode, simulating human interaction to ensure task completion.
@@
We evaluated our API-based and Hybrid Agents on WebArena, a benchmark for real-world web tasks (Zhou et al., 2023), and the results are shown in Figure 1. Our experiments revealed three key findings: (1) The API-based agent consistently outperforms browsing-based agents on WebArena tasks by around 15% on average, regardless of the comprehensiveness of APIs. (2) The API-based agent yields a higher success rate on websites with extensive API support (e.g., Gitlab) compared to those with limited API support (e.g., Reddit). This result underscores the importance of developing comprehensive API support for more accurate and efficient web task automation in the future. (3) The Hybrid Agent outperforms solely browsing-based agents and solely API-based agents, further improving accuracy by more than 5% compared to the API-based agent. By dynamically switching between approaches, the Hybrid Agent is able to provide more consistent and reliable outcomes.
@@ 2.1 The Web Browsing Task
Various benchmarks have been developed to evaluate the performance of web browsing agents. MiniWoB (Miniature World of Bits) is an early benchmark that provides simple web-based tasks such as clicking links or typing into forms, but it remains limited in complexity and realism (Shi et al., 2017). Mind2Web scales up these tasks, introducing more sophisticated interactions across websites, but it often lacks the dynamic, real-world scenarios found on the broader web (Deng et al., 2023). WebArena (Zhou et al., 2023) advances web browsing benchmarks by creating reproducible sandboxes of a variety of websites, such as managing repositories, posting online, performing online shopping, and planning trips using map services, while VisualWebArena extends WebArena to the vision modality (Koh et al., 2024a).
In this paper, we focus on WebArena tasks, which simulate real-world scenarios to evaluate an agent’s ability to complete diverse web-based activities.222Notably, upon investigation of VisualWebArena we found that APIs for handling images were relatively limited, and hence we chose to experiment on text-only tasks in this paper. Tasks in WebArena include interacting with platforms like Gitlab (to manage projects and repositories), Reddit (to browse and post content), e-commerce websites (for shopping), and mapping services (for trip planning) (Zhou et al., 2023). Task success is evaluated in three ways: (1) if the task requires producing a specific output, the agent’s response is checked for correctness; (2) for tasks involving changes to a website’s state (e.g., adding an item to a shopping cart), success is measured by verifying whether the state has changed as expected, such as ensuring the correct item and quantity have been added to the cart; and (3) if the task involves navigation, success is determined by whether the agent reaches the correct URL displaying the desired content.
@@ 2.2 A Baseline Web Browsing Agent
While there are a wide variety of agents proposed for such web navigation tasks, in this work we build upon the WebArena baseline agent (Zhou et al., 2023), which operates purely through web interaction by leveraging the accessibility tree333https://developer.mozilla.org/en-US/docs/Glossary/Accessibility_tree, a structure that exposes interactive elements like buttons, input fields, and hyperlinks (Yao et al., 2023; Gu et al., 2024). Each element of the accessibility tree is characterized by its functionality such as a hyperlink, its content, and specific web attributes (Liu et al., 2024b; He et al., 2024; Lù et al., 2024). This exposes web page elements in a hierarchical structure that is easy for agents to navigate (Samuel et al., 2024; Burns et al., 2022).
@@
While agents utilizing this method can navigate arbitrary web pages and often perform well on simpler layouts, challenges arise with the complexity of the accessibility tree. Many large language models (LLMs) are not familiar with this structure, leading to difficulties in completing tasks that require numerous or complex interactions. As a result, the average accuracy hovers in the low double digits (Liu et al., 2024a; Deng et al., 2023; Fu et al., 2024). These methods also struggle with content that need to be dynamically loaded or contents not immediately visible within the tree (Abramovich et al., 2024; Chen et al., 2024b; Lutz et al., 2024).
To give one motivating example, in Figure 2, we demonstrate a task where the agent needs to perform a task determining the number of commits made by the user SaptakS in a repository named a11yproject. Specifically, for each task, the agent is given a fixed number of steps within which it has to finish the task. Using a traditional web-browsing approach, the agent follows a complex trajectory, starting with logging into the website, navigating to the correct project, accessing the repository, and finally attempting to view the list of commits. However, due to the large number of commits made by other users, the commits by SaptakS are located much further down on the web page, requiring the agent to scroll down many times. As a result, despite completing 15 actions, the browsing agent is unable to retrieve the required information.
@@ 3 From Web Browsing to API Calling
In contrast to web browsing, API calling offer a direct interface for machines to communicate with web services, reducing operational complexity. In this section, we explore an API-based approach when performing web tasks.
@@ 3.1 APIs and API Documentation
For websites that offer API support, pre-defined endpoints can be utilized to perform tasks efficiently. These APIs, following standardized protocols like REST444https://en.wikipedia.org/wiki/REST, allow interaction with web services through sending HTTP requests (e.g., GET, POST, PUT) and receiving structured data such as JSON objects555https://www.json.org/json-en.html as responses. Websites often provide official documentation for the APIs, which can give guidance on how to utilize the APIs. Some documentation is provided in README 666https://en.wikipedia.org/wiki/README format, some are in OpenAPI YAML777https://yaml.org/ format, and some are in plain text format. For instance, Figure 3 shows the official README documentation of a Gitlab API GET /api/{id}/commits. It documents the functionality, input arguments, and output types of the API. For example, one could use the Python requests library, by calling requests.get("gitlab.com/api/a11yproject/commits"), to retrieve all commits of the repository a11yproject. This would return a JSON list containing all the commits to this repo, as shown in Figure 3.
@@ 3.2 Obtaining APIs for Agents
There are some challenging situations where APIs exist but good documentation is not officially available. In such cases, additional steps may be required to obtain a list of accessible APIs. In this case, we inspected the frontend or backend code of the website to extract undocumented API calls that can still be utilized by the agent. Then, based on the implementation of APIs of the website, leverage an LLM (GPT-4o888https://openai.com/index/hello-gpt-4o/) to generate these YAML or README files. By prompting GPT-4o with the relevant implementation details of the APIs (for example, the implementation files of the APIs or example traces of API calls), we generate comprehensive documentation, including input parameters, expected outputs, and example API calls.
@@ 3.3 Using APIs in Agents
For websites with a smaller number of API endpoints999Specifically, we use a threshold of 100 APIs, but this could be adjusted depending on the supported language model context size., we directly incorporate the full documentation into the prompt provided to the agent. This approach of directly feeding the full documentation worked well for websites with a limited number of API endpoints, as it allowed the agent to have immediate access to all the necessary information without the need for a more complex retrieval mechanism.
@@
For websites with a larger number of endpoints, providing the full documentation directly within the prompt was impractical due to the size limitations of agent inputs. To address this, we employ a two-stage documentation retrieval process, which allowed access to only the relevant information as needed, keeping the initial prompt concise.
In the first stage, the user prompt provide a description of the task, with a list of all available API endpoints along with a very brief description of each API. For example, {"GET /api/{id}/commits": "Get a list of commits in a project"}. This initial summary helps facilitating understanding the scope of all the available APIs while staying within the prompt size constraints.
In the second stage, if the model determines that it needs detailed information about a specific API endpoint or some API endpoints, it can use a tool called get_api_documentation. This tool maintains a dictionary that maps each API to its API documentation respectively. The dictionary is generated using pattern match in Python to retrieve substrings related to each endpoints. get_api_documentation is able to searche the dictionary and retrieve the full README or YAML documentation for any given endpoint by calling get_api_documentation with the endpoint’s identifier. This might include the input parameters, output formats, and examples of how to interact with the endpoint. For example, to retrieve the documentation for the endpoint GET /api/id/commits, the agent would call get_api_documentation("GET /api/id/commits"), and an example returned API documentation is the documentation in Figure 3.
This retrieval method allows the agent to make flexible and informed choices during the execution of tasks. If the agent finds that an API does not meet its needs or if it encounters an error, it can easily retrieve the documentation for a different API endpoint by calling the function again. This dynamic approach promotes adaptability and minimizes the risk of incorrect API usage when the number of APIs available is large. The prompt can be found in Appendix A.2.
@@ 4 Hybrid Browsing+API Calling Agents
We have proposed API-based methods for handling web tasks, but the question arises: given the benefits of API calling, should we discard web browsing altogether? The most obvious bottleneck is that not all websites offer comprehensive API support. Some platforms offer limited or poorly documented APIs (e.g. there is no API for shopping on Amazon101010https://www.amazon.com), forcing agents to rely on traditional web browsing methods to complete tasks.
To deal with these situations, we propose a hybrid methods that integrates both browsing-based and API-based approaches, and developed a Hybrid Agent capable of interleaving API calls and web browsing, switching dynamically based on task requirements and the available resources. Specifically, for each task, the agent is given the fixed step budget within which it has to finish the task. In each step, the agent could either (1) communicate with humans in natural language to ask for clarification or confirmation, or 2) generate and executes Python code which could include performing API calling, or 3) performs web browsing actions. The agent could choose freely among these three options, depending on the agent’s confidence which method could best tackle the task.
The ideal case is that for websites that offer comprehensive API support, the Hybrid Agent can utilize well-documented endpoints to perform tasks more efficiently than it could through web browsing; for websites with limited API support or poorly documented APIs, the Hybrid Agent could rely more on web browsing to fulfill certain tasks. We later find that enabling an agent to interleave API calling and web browsing boost the agent’s performance (see Section 6).
@@ 5.1 Dataset Description
For our experiments, we utilized the WebArena dataset (Zhou et al., 2023) as the primary evaluation benchmark. WebArena is a comprehensive benchmark designed for real-world web tasks, providing a diverse set of websites that simulate various online interactions. The tasks within WebArena reflect common user activities such as navigating websites, performing administrative tasks, and posting online.
The dataset mainly includes five distinct websites, each containing various intents representing different tasks: Gitlab, Map, Shopping, Shopping Admin, Reddit, and Multi-Website Tasks. We include a more detailed descriptions of the tasks in Appendix A.1. This diverse set of websites and tasks within WebArena allows for a comprehensive evaluation of the agents, testing their ability to handle both API-based interactions and web browsing across varied web settings.

### v2

@@ Abstract
Web browsers are a portal to the internet, where much of human activity is undertaken. Thus, there has been significant research work in AI agents that interact with the internet through web browsing. However, there is also another interface designed specifically for machine interaction with online content: application programming interfaces (APIs). In this paper we ask – what if we were to take tasks traditionally tackled by Browsing Agents, and give AI agents access to APIs? To do so, we propose two varieties of agents: (1) an API-calling agent that attempts to perform online tasks through APIs only, similar to traditional coding agents, and (2) a Hybrid Agent that can interact with online data through both web browsing and APIs. In experiments on WebArena, a widely-used and realistic benchmark for web navigation tasks, we find that API-Based Agents outperform web Browsing Agents. Hybrid Agents out-perform both others nearly uniformly across tasks, resulting in a more than 24.0% absolute improvement over web browsing alone, achieving a success rate of 38.9%, the SOTA performance among task-agnostic agents. These results strongly suggest that when APIs are available, they present an attractive alternative to relying on web browsing alone.
[…]
Beyond Browsing: API-Based Web Agents
Yueqi Song, Frank Xu, Shuyan Zhou, Graham Neubig {yueqis,gneubig}@cs.cmu.edu Carnegie Mellon University
@@ 1 Introduction
Web agents use browsers as an interface to facilitate humans in performing daily tasks such as online shopping, online planning, trip planning, and other work-related tasks (Liu et al., 2018; Li et al., 2020; Rawles et al., 2023; Patil et al., 2023; Pan et al., 2024; Chen et al., 2024a; Huang et al., 2024; Durante et al., 2024). Existing web agents typically operate within the space of graphical user interfaces (GUI) (Zhang et al., 2023; Zhou et al., 2024; Zheng et al., 2024), using action spaces that simulate human-like keyboard and mouse operations, such as clicking and typing. To observe webpages, common approaches include using accessibility trees, a simplified version of the HTML DOM tree, as input to text-based models (Zhou et al., 2024; Drouin et al., 2024a), or multi-modal, screenshot-based models (Koh et al., 2024a; Xie et al., 2024; You et al., 2024; Hong et al., 2023). However, regardless of the interaction method with websites, there is no getting around the fact that these sites were originally designed for humans, and may not be the ideal interface for machines.
@@
Notably, there is another interface designed specifically for machine interaction with the web: application programming interfaces (APIs) (Chan et al., 2024). APIs allow machines to communicate directly with backends of web services (Branavan et al., 2009), sending and receiving data in machine-friendly formats such as JSON or XML (Meng et al., 2018; Xu et al., 2021; Trivedi et al., 2024). Nonetheless, whether AI agents can effectively use APIs to tackle real-world online tasks, and the conditions under which this is possible, remain unstudied. In this work, we explore methods for tackling tasks normally framed as web-navigation tasks with an expanded action space to interact with APIs. To do so, we develop new API-Based Agents that directly interact with web services via API calls. This method bypasses the need to interact with web GUIs.
@@
At the same time, not all websites have extensive API support, in which case web browsing actions may still be required. To address these cases, we explore a hybrid approach that combines API-Based Agents with Browsing Agents, as described in Figure 1. By implementing an agent capable of interleaving API calls and web browsing, we found that agents benefit from the flexibility of this hybrid model. When APIs are available and well-documented, the agent can directly interact with the web services. For websites with limited API support, the agent seamlessly interleaves API calling and browsing to ensure task completion.
@@
We evaluated our API-Based and Hybrid Agents on WebArena, a benchmark for real-world web tasks (Zhou et al., 2024), and the results are shown in Figure 1. Our experiments revealed three key findings: (1) The API-Based Agent outperforms the Browsing Agent on WebArena by around 15% on average. (2) The API-Based Agent yields a higher success rate on websites with good API support (e.g., Gitlab) compared to those with limited API support (e.g., Reddit). This result underscores the importance of developing comprehensive API support for more accurate and efficient web task automation in the future. (3) The Hybrid Agent outperforms solely Browsing and solely API-Based Agents, further improving accuracy by 6% compared to the API-Based Agent. By dynamically interleaving approaches, the Hybrid Agent is able to provide more consistent and reliable outcomes.
@@ 2.1 The Web Browsing Task
Various benchmarks have been developed to evaluate web Browsing Agents. MiniWoB (Miniature World of Bits) is an early benchmark that provides simple web-based tasks such as clicking links or typing into forms, but it remains limited in complexity and realism (Shi et al., 2017). Mind2Web scales up these tasks, introducing more complex interactions across websites, but it primarily focuses on basic web operations (Deng et al., 2023). WebArena (Zhou et al., 2024) advances web browsing benchmarks by creating reproducible sandboxes of various websites, such as managing repositories, posting online, performing online shopping, and planning trips using map services, while VisualWebArena extends WebArena to the vision modality (Koh et al., 2024a).
In this paper, we focus on WebArena tasks, which simulate real-world scenarios to evaluate an agent’s ability to complete diverse web-based activities.111Notably, upon investigation of VisualWebArena we found that APIs for handling images were relatively limited, and hence we chose to experiment on text-only tasks in this paper. WebArena tasks include interacting with platforms like Gitlab (to manage projects and repositories), Reddit (to browse and post content), e-commerce websites (for shopping), and mapping services (for trip planning) (Zhou et al., 2024). Task success is evaluated in three ways: (1) if the task requires producing specific outputs, agents’ responses are checked for correctness; (2) for tasks involving changes to a website’s state (e.g., adding items to shopping carts), success is verified by whether the state has changed as expected, such as ensuring the correct items have been added to the cart; and (3) if the task involves navigation, success is determined by whether the agent reaches the correct URL displaying the desired content.
@@ 2.2 A Baseline Web Browsing Agent
While there are a wide variety of agents proposed for such web navigation tasks, in this work we build upon the WebArena baseline agent (Zhou et al., 2024), which operates purely through web interaction by leveraging the accessibility tree222https://developer.mozilla.org/en-US/docs/Glossary/Accessibility_tree, a structure that exposes interactive elements like buttons, input fields, and hyperlinks (Yao et al., 2023; Gu et al., 2024). Each element of the accessibility tree is characterized by its functionality such as a hyperlink, its content, and specific web attributes (Liu et al., 2024b; He et al., 2024a; Lù et al., 2024). This exposes webpage elements in a hierarchical structure that is easy for agents to navigate (Samuel et al., 2024; Burns et al., 2022).
@@
While agents utilizing this method can navigate arbitrary webpages and often perform well on simple layouts, challenges arise with the complexity of GUIs. Many large language models (LLMs) are not familiar with accessibility trees, which leads to difficulties in completing tasks that require numerous or complex interactions, resulting in lower accuracies (Liu et al., 2024a; Deng et al., 2023; Fu et al., 2024). These methods also struggle with content that need to be dynamically loaded or contents not immediately visible within the tree (Abramovich et al., 2024; Chen et al., 2024b; Lutz et al., 2024).
To give a motivating example, in Figure 2, we demonstrate a task where agents need to determine the number of commits made by the user SaptakS in a repository named a11yproject. For each task, agents are given a fixed number of steps within which to complete the task. Using a traditional browsing approach, the agent follows a complex trajectory, starting with logging in, navigating to the correct project, accessing the repository, and finally attempting to view the list of commits. However, due to the large number of commits made by other users, the commits by SaptakS are located much further down on the webpage, requiring agents to scroll down many times. As a result, despite completing 15 steps, the Browsing Agent is unable to retrieve the required information.
@@ 3 From Web Browsing to API Calling
In contrast, API calling allows machines to directly communicate with web services, reducing operational complexity. In this section, we explore an API-based approach when performing web tasks.
@@ 3.1 APIs and API Documentation
For websites that offer API support, pre-defined endpoints can be utilized to perform tasks efficiently. These APIs, following standardized protocols like REST333https://en.wikipedia.org/wiki/REST, allow interaction with web services through sending HTTP requests (e.g., GET, POST, PUT) and receiving structured data such as JSON objects444https://www.json.org/json-en.html as responses. Websites often provide official documentation for the APIs, which can give guidance on how to utilize the APIs. Some documentation is provided as plain text, some in README 555https://en.wikipedia.org/wiki/README format, and some in OpenAPI YAML666https://yaml.org/ format. Figure 3 shows an example of the Gitlab README documentation of GET /api/{id}/commits, documenting it’s functionality, input arguments, and output types. For example, to retrieve all commits to a11yproject, one could use the Python requests library, by calling requests.get“gitlab.com/api/a11yproject/commits”). This returns a JSON list containing all the commits to this repo, as shown in Figure 3.
@@ 3.2 Obtaining APIs for Agents
There are some challenging situations where APIs exist but good documentation is not officially available. In such cases, additional steps may be required to obtain a list of accessible APIs. In this case, we inspected the frontend or backend code of the website to extract undocumented API calls that can still be utilized by the agent. Then, based on the implementation of APIs of the website, leverage an LLM (GPT-4o777https://openai.com/index/hello-gpt-4o/) to generate these YAML or README files. By prompting GPT-4o with the relevant implementation details of the APIs (for example, the implementation files of the APIs or example traces of API calls), we generate comprehensive documentation, including input parameters, expected outputs, and example API calls.
@@ 3.3 Using APIs in Agents
For websites with smaller numbers of APIs888We use a threshold of 100 APIs, but this could be adjusted depending on the supported language model context size., we directly incorporate the full documentation into the prompt provided to the agent. This approach of directly feeding the full documentation worked well for websites with a limited number of API endpoints, as it allowed the agent to have immediate access to all the necessary information without the need for a more complex retrieval mechanism.
@@
For websites with more APIs, providing the full documentation in the prompt is impractical due to size limitation of agent inputs. To address this, we use a two-stage documentation retrieval process, allowing access to only the needed information to keep the initial prompt concise.
In the first stage, the user prompt provides a task description, with a list of all available APIs along with a brief description of each. For example, {“GET /api/{id}/commits”: “List commits in a project”}. This initial summary helps in understanding the scope of all the available APIs while staying within the prompt size constraints.
In the second stage, if the model determines that it needs detailed information about one or more specific API endpoints, it can use a tool named get_api_documentation, which maintains a dictionary that maps each API to its documentation respectively. The dictionary is generated using Python pattern match to retrieve substrings related to each endpoints. This tool is able to search the dictionary and retrieve the full README or YAML documentation for any given endpoint with the endpoint’s identifier. The resulting documentation might include the input parameters, output formats, and examples of how to interact with the endpoint. For example, to retrieve the documentation for the API GET /api/id/commits, the agent would call get_api_documentation(‘‘GET /api/id/commits’’). An example returned API documentation is the documentation in Figure 3.
This retrieval method allows the agent to make flexible and informed decisions to perform tasks. If the agent finds that an API does not meet its needs or if it encounters an error, it can easily retrieve the documentation for a different API by calling the tool again. This dynamic approach promotes adaptability and minimizes the risk of incorrect API usage when the number of APIs available is large. The prompt can be found in Appendix A.6.
@@ 4 Hybrid Browsing+API Calling Agents
We have proposed API-based methods for handling web tasks, but the question arises: given the benefits of API calling, should we discard browsing altogether? The most obvious bottleneck is that not all websites offer good API support. Some platforms offer limited or poorly documented APIs (e.g. no API for shopping on Amazon999https://www.amazon.com), forcing agents to rely on browsing to complete tasks.
To deal with these situations, we propose a hybrid methods that integrates both browsing and API calling, and developed a Hybrid Agent capable of dynamically interleaving API calls and web browsing based on task requirements and the available resources. Specifically, for each task, the agent is given the fixed step budget within which it has to finish the task. In each step of a task, the agent could either (1) communicate with humans in natural language to ask for clarification, or 2) generate and executes Python code which could include performing API calling, or 3) performs web browsing actions. The Hybrid Agent could choose freely among these options, depending on the agent’s confidence in which method is the best for each step.
Ideally, for websites with good API support, the Hybrid Agent can utilize well-documented APIs to perform tasks more efficiently than it could through only browsing; for websites with limited API support or poor documentation, the Hybrid Agent could rely more on browsing. We find that enabling it to interleave API calling and web browsing boosts task performance (see Section 6).
@@ 5.1 Dataset Description
We utilized WebArena (Zhou et al., 2024) as the primary evaluation benchmark. WebArena is a comprehensive benchmark designed for real-world web tasks, providing a diverse set of websites that simulate various online interactions, allowing comprehensive evaluation of agents’ abilities to handle both API calling and web browsing across varied web settings. WebArena mainly includes five websites, each with various intents representing different tasks: Gitlab, Map, Shopping, Shopping Admin, Reddit, and Multi-Site tasks. A detailed descriptions of the tasks is in Appendix A.2.